answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

How many distinct arrangements can be made with the letters in the word BOXING?

The number of distinct arrangements of the letters of the word BOXING is the same as the number of permutations of 6 things taken 6 at a time. This is 6 factorial, which is 720. Since there are no duplicated letters in the word, there is no need to divide by any factor.

What Percent of data is below the mean in a normal distribution?

In the normal distribution, the mean and median coincide,

and 50% of the data are below the mean.

What the size of the sample space for rolling a fair die?

When a fair die is rolled, there are 6 possible outcomes {1,2,3,4,5,6}.

The sample space consists of 6 points, so its size is 6.

What is the z score for 96 percent?

1.75 using table for standard normal cumulative probabilities

What are two characteristics of semen that enhance the probability of fertilization?

The first is the concentration of sperm cells in semen - the higher the concentration better the possibility of conception. The second is the percentage of normal sperm cells - the lesser the number of abnormal sperm cells, the better the possibility of conception.

What is the expected shape of the distribution of the sample mean?

The distribution of the sample mean is bell-shaped or is a normal distribution.

What is the probability of drawing a heart and rolling a 4?

The probability of drawing a heart in a single random draw from a normal deck of playing cards and rolling a 4 on a fair number cube is:

1/4*1/6 = 1/24

What is the probability that there will be rain during at least one of three days?

There is insufficient information in the question to properly answer it. You need to specify what range of dates and what the typical climate is like, perhaps by specifying the geographic location that you are interested in.

If you want to take this to an extreme, and say for any date and for anywhere on the world, then since it is always raining at least somewhere, then the probability that there will be rain during at least one of three days is 1.

Is it possible to doubt everything or almost everything?

It's often less difficult and requires less effort to doubt rather than to believe.

It requires less thinking, rationalization, and research to doubt a subject than it

would to do the necessary steps to bring one to raw belief.

In order to doubt, only a few factors must be considered. However, in order to reach pure belief, many factors, theorems, and proofs must be presented.

And, still, if one is leaning toward non-belief, it is difficult to separate the

salient points that led you not to believe against the other side of the debate that would instill in you a true belief.

What does it mean if you have the same birthday as someone else?

It simply means that you were born on the same day of the year.

From a probabilistic standpoint, assuming a uniform distribution, you have a 1 in 365.25 chance of matching a specific someone else's birthday.

With nearly 7 billion people on this planet, there are probably about 19 million people with each of the possible birthdays living.

How many ways can you arrange 5 books on a shelf order is important?

5 books can be lined up on a shelf in (5 x 4 x 3 x 2 x 1) = 120 different sequences.

Seven coins are tossed What is the probability of 3 heads or 2 tails?

The probability of one event or the other occurring is the probability of one plus the probability of the other. The probability of getting 3 heads is the probability of 3 heads (1/23) multiplied by the probability of 4 tails (1/24) multiplied by the number of possible ways this could happen. This is 7c3 or 35. Thus the probability of 3 heads is 0.2734375. The probability of 2 tails is the probability of 2 tails (1/22) multiplied by the probability of 5 heads (1/25) multiplied by the number of ways this could happen. That is 7c5 or 21. Thus the probability of 2 tails is 0.1640625

The probability of one or the other is the sum of their probabilities:

0.1640625 + 0.2734375 = 0.4375

Thus the probability of getting 3 heads or 2 tails is 0.4375.

What is the shape of the distribution of the mean of a sample?

The mean of a sample is a single value and so its distribution is a single value with probability 1.

What is the probability of getting a diamond and a black 7?

The probability of getting a diamond and a black seven is zero. Diamonds are red.

What is the probability of of tossing heads on the first 6 tosses of a fair coin?

The probability of tossing heads on all of the first six tosses of a fair coin is 0.56, or 0.015625.

The probability of tossing heads on at least one of the first six tosses of a fair coin is 1 - 0.56, or 0.984375.

Ap statistics multiple choice preparation book answers?

Reddish peach book (authors: Joyce Heller, Lenore Lerer, and others) answers

Section 1 sample examination 1

1. For the years 1950-1980… A

2. A

3. C

4. D

5. B

6. C

7. B

8. E

9. C

10. E

11. E

12. D

13. B

14. C

15. E

16. D

17. D

18. B

19. B

20. E

21. B

22. C

23. C

24. B

25. D

26. E

27. C

28. B

29. B

30. C

31. D

32. C

33. E

34. D

35. B

36. C

37. D

38. B

39. C

40. A student was interested in the relationship… D

Sample examination 2

1. Creating a sample of students by starting… D

2. D

3. B

4. A

5. B

6. B

7. B

8. B

9. C

10. E

11. E

12. A

13. A

14. D

15. E

16. E

17. D

18. B

19. B

20. D

21. D

22. D

23. A

24. B

25. C

26. B

27. C

28. C

29. E

30. E

31. D

32. B

33. C

34. D

35. D

36. D

37. A

38. C

39. D

40. C

What is the probability of getting a sum greater than 9 when two dice are rolled?

There are 30 possible combinations of die rolls (assuming 2 6-sided die) (excluding "doubles")

The highest total you can get is 12 so the numbers you are after are 10,11 and 12...

The possible combos which add to greater than 9 are:

4,6

6,4

5,5

5,6

6,5

So the probability of rolling a 10, 11 or 12 is 5:30 (5 times out of 30 rolls)

What are the odds of running out of peanut butter and jelly at the same time?

That's only possible to calculate if more detail is given. For example, how likely is the person to buy jelly normally? How much of it do they consume on average? how likely are they to buy peanut butter? what is their average consumption? Do they actually enjoy eating peanut butter and jelly at the same time? etc...

What is 25 percent of 132?

Oh, dude, math time! So, like, to find 25 percent of 132, you just multiply 132 by 0.25. That gives you 33. But like, who even needs to know that? Just use a calculator or something, man.